OpenStudy (anonymous):

5 sqrt(18)+4 sqrt(32)

OpenStudy (anonymous):

5√18+4√32

OpenStudy (anonymous):

which is 31sqrt2

OpenStudy (anonymous):

yea i checked it on wolfram alpha but wen i did it i ended up with 23 sqrt2

OpenStudy (anonymous):

i hoping for a step by step guide or som tyoe of aid

OpenStudy (anonymous):

√18 can be written as √(9*2). The √ is distributive over multiplication, so √(9*2) = √9 * √2 = 3√2. Same thing with √32: √32 = √(16*2) = √(8*2*2) = √(4*4*2) = √4√4√2 = 2 * 2 √2 = 4√2. So, 5√18 + 4√32 = 5 (3√2) + 4 (4√2) = 15√2 + 16√2 = 31√2

OpenStudy (anonymous):

Oops, the quicker way to simplify √32 would be recognize the √16 is 4 immediately. I overlooked that and decided to simplify it further into 2's haha.

OpenStudy (anonymous):

thank you for ur help it was very helpful